3H↓2 + N↓2 → 2NH↓3
calculate the minimum volume of hydrogen required to completely convert 1000dm^3 of nitrogen into ammonia
1
Expert's answer
2018-09-10T08:01:28-0400
V(N2) = 1000 dm3
n (N2) = V (N2)/Vm=1000/22.4=44.643 mole
n(H2) = 44.643 * 3 =133.93 mole (3 is stoichiometric coefficient before hydrogen in the reaction equation)
V(H2)=n(H2)*Vm=3000 dm3